### v1

@@ Abstract
Index Terms:
cryptography, prime factorization, public key, quantum computing, Shor’s algorithm
I Introduction
Quantum computers have the potential to transform computing by efficiently solving certain problems that are intractable for classical machines. Shor’s algorithm [b1, b2] is particularly notable for its ability to factor large integers in polynomial time, outperforming the best-known classical methods. At the core of Shor’s algorithm is the Quantum Fourier Transform (QFT), which enables key subroutines to be implemented exponentially faster than their classical counterparts. This capability threatens widely used public-key cryptographic protocols, such as RSA and also ECC [b29], whose security is based on the presumed hardness of integer factorization or related number-theoretic problems. While theoretical advancements have clarified how quantum computers could eventually compromise such systems, the practical implementation of Shor’s algorithm on existing quantum hardware remains extremely challenging — even for integers far smaller than those used in deployed cryptographic schemes. In the current noisy intermediate-scale quantum (NISQ) [preskill_quantum_2018] era, quantum devices have limited qubit counts, significant noise, and short coherence times. Understanding what they can and cannot realistically achieve is crucial for assessing the urgency of quantum-safe cryptographic migration.
@@ I Introduction
II Understanding Different Quantum Computers
Quantum computing is built on qubits as the fundamental units of quantum information. Broadly, qubit implementations can be categorized into two families: synthetic qubits, which are artificially engineered entities designed to effectively realize two-level quantum behavior, and natural qubits, which directly use the intrinsic quantum states of atoms, ions, or photons. Table I provides an overview of various key qubit technologies, representative commercial providers, and high-level performance characteristics.
TABLE I: Quantum computing with different qubit technologies and their characteristics.
@@ II Understanding Different Quantum Computers
II-A Quantum Computers with Synthetic Qubits
@@
II-B Quantum Computers with Natural Qubits
@@
III Review on Shor’s Implementations
III-A Shor’s Algorithm
As discussed in [b18], quantum computers excel at solving problems with modest classical input/output data but high computational complexity. Shor’s algorithm [b1, b2] is a canonical example: it factors an $N$-bit integer in time polynomial in $N$, in stark contrast to the sub-exponential but super-polynomial complexity of the best-known classical factoring algorithms.
@@ III-A Shor’s Algorithm
III-B Shor’s Implementation: State-of-the-Art
@@
In 2001, a group of IBM researchers presented the earliest experimental demonstration of Shor’s algorithm, successfully factoring $15=3\times 5$ using a seven-qubit liquid-state  [^1] Uses nuclear spins in molecules dissolved in a liquid as qubits. nuclear magnetic resonance (NMR) quantum computer [b3]. While liquid-state NMR quantum computers are not considered a scalable route for solving larger problems, the experiment provided an important proof of principle, demonstrating the algorithm’s feasibility.
@@
Subsequent experiments implemented compiled versions of Shor’s algorithm using photonic qubits, again targeting the factorization of 15 [b4, b5]. These demonstrations focused on generating and characterizing multi-qubit entanglement in circuits derived from Shor’s algorithm, but relied on heavy compilation that exploits prior knowledge of the factors.
@@
In 2012, factorization of 15 was successfully demonstrated on a superconducting quantum processor [b6]. This milestone marked a significant advancement, showcasing the feasibility of implementing Shor’s algorithm on a quantum platform aligned with long-term scalability goals. In the same year, a photonic implmentation successfully factored 21 [b7].
@@
In 2016, factorization of 15 was achieved using a trapped-ion implementation which used a qubit recycling technique [b8]. Unlike previous compiled approaches, this demonstration did not rely on prior knowledge of the prime factors, and the design was argued to be scalable in principle, albeit still limited to a four-bit composite.
@@
More recent work was carried out on larger superconducting devices. In 2019, an attempt was made to factor 35 on the IBM Q System One [b9]. However, the algorithm achieved only a 14% success rate due to accumulated noise, illustrating the fragility of deep circuits on current hardware. The latest IBM-based demonstation adopts a hybrid quantum-classical technique to variationally factor 253 (an 8 bit composite) [b30]. While such methods are promising, they change the algorithmic structure and require careful interpretation when extrapolating to cryptographic scales.
@@
A number of works claim factorizations of larger integers using Shor-like circuits. As emphasized in the aptly-titled manuscript “Oversimplifying Quantum Factoring” [b12], many of these experiments rely on heavily compiled circuits that incorporate knowledge of the answer, drastically lowering resource requirements, and sometimes reducing the problem to something closer to classical coin flipping than genuine quantum factoring. This underscores the importance of benchmarking implementations that preserve the true complexity of order-finding.
@@
In parallel, classical simulations of Shor’s algorithm on large-scale high-performance computing systems have advanced significantly. Recent work [b19] reached 39 bits RSA key length on GPU clusters, validating the algorithmic structure and resource trends in a controlled environment. Nevertheless, the current (public) record is still held by classical algorithms such as the “Generalized Number Field-Sieve” (GNFS) which has reached 829 bits of key length [b25].
@@
On the algorithmic side, the past two years have seen the first notable developments in the 30 years since Shor’s discovery in 1994. Recent work has lowered qubit counts [b20, b21, b26] and improved asymptotic runtime [b22], particularly by leveraging advances in quantum error correction and more efficient order-finding techniques. An important point to observe is the growing efficiency of incorporating progress in quantum error correction, which can be considered as the software layer between hardware (noisy physical qubits) and application (factoring). For example, updated resource estimates for factoring RSA-2048 have improved from around 20 million qubits and 8 hours of execution time [b17] to fewer than one million qubits and roughly 5 days [b27], under realistic assumptions about error rates and QEC overhead. Architectures that combine novel qubit designs with multimode memories could, at least in principle, reduce the required qubit count by another two orders of magnitude, down to approximately 13 436 qubits [b28]. These developments emphasize that both hardware and algorithms are progressing and that resource estimates continue to evolve.
IV Shor’s Implementation on Existing Quantum Platforms
IV-A Setup and Notation
@@
However, with modern post-processing (Ekerå et al. [b31]), the required phase precision can be cut to roughly $t\approx n+O(\log n)$ while maintaining high single-run success via limited classical searches, thereby reducing depth and repetitions [b31, b32]. Thus, we allow ourselves to use a lower $t=10<12$ for the case $N=35$ while still considering the framework scalable  [^2] Further details on the experimental certification method and refinements regarding scalability in the context of Shor’s algorithm are provided in the Appendix A-A..
IV-B Public Quantum Platforms
@@
TABLE II: Specifications of the ibm_torino Quantum Processor
@@
Table II summarizes the characteristics of ibm_torino, the backend used in our experiments. Two coherence-time metrics are listed. $T_{1}$ (energy relaxation time) measures how long a qubit prepared in the excited state ($|{1}\rangle$) remains there before decaying to the ground state ($|{0}\rangle$). Longer $T_{1}$ values indicate reduced energy relaxation and better stability. $T_{2}$ (dephasing time) quantifies how long a qubit maintains phase coherence in a superposition state; it captures the rate at which quantum information is lost due to environmental noise and control imperfections. While $T_{2}$ is upper-bounded by $2\,T_{1}$, it is typically significantly shorter in practice due to imperfect control and environmental noise.
@@
Our implementation is based on standard constructions for reversible arithmetic. We employ a Cuccaro-style ripple-carry adder [Cuccaro2004] and a two’s-complement overflow/comparator method for modular addition [Oumarou2022], avoiding QFT-space adders [Draper2000]. Controlled powers of $a$ are implemented via precomputed exponents $a^{2^{k}}\bmod N$ within a textbook QPE order-finding scaffold [Kitaev1997, Cleve1998] and we target Qiskit Runtime Sampler V2 [QiskitRuntime].
@@
We implement the parallel (non-iterative) phase-estimation variant of Shor’s order-finding rather than Kitaev’s iterative version [Kitaev1997]. Both are algorithmically equivalent: the dominant cost is the controlled modular exponentiation $U_{a}$, which scales as $\tilde{O}(n^{3})$ gates for $n=\lceil\log_{2}N\rceil$ with standard reversible adders and modular reduction, while the inverse QFT contributes $O(t^{2})$ gates for $t$ phase bits. The choice is therefore an engineering trade-off. Parallel QPE uses $t$ control qubits and a single inverse QFT, yielding a full $t$-bit outcome per shot. Iterative QPE reuses one control qubit over $t$ rounds with mid-circuit measurement and classical feed-forward, thereby reducing qubit count but increasing depth adding additional SPAM/latency errors. In our experiments, $t\approx 2\lceil\log_{2}N\rceil$ is modest, so the extra control qubits are acceptable and let us (i) avoid real-time feedback, (ii) exploit vectorized sampling, and (iii) apply a uniform multi-peak histogram analysis pipeline based on acceptance windows and binomial tests. In a fault-tolerant setting, either variant remains viable; our choice is driven by hardware constraints, not complexity.
@@ IV-C Results (N=15, 21, 35)
We then executed the same circuits on ibm_torino, after transpilation to the device topology. For each modulus $N$, we focus on detecting a statistically significant excess of probability mass in the theoretically predicted windows around $sL/r$ relative to the uniform baseline. The formulas of the uniform baseline and the empirical probability used for the experiments, rigorously derived in the Appendix A-A, are the following:
@@
We consider three composite numbers: $N=15$, $N=21$, and $N=35$. In the Appendix subsection A-G, we further discuss the shape of the empirical histograms compared to the ideal noiseless expected distributions.
@@ V Limitations and Challenges
To achieve genericity at fixed width, the arithmetic would need to treat $N$ and related parameters as data rather than as compile-time structure. This would entail allocating an $n$-qubit modulus register $\mathrm{NREG}$ (and addend/multiplier registers as needed), replace constant-injection adders with register–register Cuccaro adders plus a reversible subtract-and-borrow comparator against $\mathrm{NREG}$, and implement modular multiplication as a variable shift-and-add with reductions modulo $\mathrm{NREG}$ (optionally via parameterized Fourier-space adders) [b11]. Consequently, a comprehensive benchmark should employ $N$-agnostic arithmetic register–register add/compare/multiply mod $N$ or parameterized Fourier-space rotations so that a single width-fixed topology works for any $N<2^{n}$. Per-$N$ specialized circuits primarily demonstrate the behavior of bespoke instances and do not fully probe a device’s readiness for the generic operations required at cryptographically relevant key sizes.
@@
Quantum machine fidelity captures how accurately a device implements its nominal operations. Other factors such as qubit connectivity, coherence times, and readout error rates provide further metrics to estimate whether reliable results are produced when running Shor’s algorithm on real hardware. While we had access to calibration data for ibm_torino (Table II), these values represent median performance and can vary substantially over time. In practice, we observed non-negligible day-to-day fluctuations that occasionally degraded circuit behavior to the point of rendering runs effectively unusable. As a result, some experiments had to be discarded as outliers corresponding to unfavorably calibrated device states. Below are presented the statistics of some error metrics over the months preceding and succeeding the experiments, for qubit q_0. A more systematic characterization of worst-case behavior—for instance, via repeated sampling over many calibration cycles—would be necessary to rigorously assess reliability guarantees. However, such an investigation would require significantly more QPU access time and lies beyond the scope of this study.
TABLE III: Calibration statistics for ibm_torino machine over the months preceding and succeeding the experiments, for qubit $q_{0}$
@@
Our experiments focus primarily on detecting a statistically significant quantum signal in QPE histograms, rather than fully establishing polylogarithmic scaling of the expected runtime for a fixed acceptance parameter $k$ (see Appendix A-F). For each small modulus $N$, we tune window sizes and shot counts to test whether the mass in acceptance regions exceeds a uniform baseline. This is sufficient to detect coherence in the presence of noise but does not yet constitute a full scalability study.
@@
VI Trends in Quantum Computing Developments
Investment in quantum technologies has surged in recent years. Global public and private funding has been estimated to exceed \$55 billion [^7] https://www.qureca.com/quantum-initiatives-worldwide/, spanning quantum computing, communication, sensing, and related areas. Many governments have launched national or regional initiatives, such as the U.S. National Quantum Initiative [^8] https://www.quantum.gov/, the European Quantum Flagship  [^9] https://qt.eu/ and major programs in China and other countries, aimed at securing leadership in this emerging and potentially transformative field [b10]. In parallel, private sector funding has grown rapidly, with large technology companies such as IBM, Google, Microsoft, and Amazon heavily investing in quantum computing, alongside startups such as IonQ, Quantinuum, Rigetti, PsiQuantum, Xanadu, and others, which are developing specialized quantum systems. Table IV illustrates selected quantum computing roadmaps announced by industrial players, grouped by qubit technology.

### v2

@@ Abstract
Keywords: cryptography, prime factorization, public key, quantum computing, Shor’s algorithm
1 Introduction
Quantum computers have the potential to transform computing by efficiently solving certain problems that are intractable for classical machines. Shor’s algorithm [1, 2] is particularly notable for its ability to factor large integers in polynomial time, outperforming the best-known classical methods. At the core of Shor’s algorithm is the Quantum Fourier Transform (QFT), which enables key subroutines to be implemented exponentially faster than their classical counterparts. This capability threatens widely used public-key cryptographic protocols, such as RSA and also ECC [3], whose security is based on the presumed hardness of integer factorization or related number-theoretic problems. While theoretical advancements have clarified how quantum computers could eventually compromise such systems, the practical implementation of Shor’s algorithm on existing quantum hardware remains extremely challenging — even for integers far smaller than those used in deployed cryptographic schemes. In the current noisy intermediate-scale quantum (NISQ) [4] era, quantum devices have limited qubit counts, significant noise, and short coherence times. Understanding what they can and cannot realistically achieve is crucial for assessing the urgency of quantum-safe cryptographic migration.
@@ 1 Introduction
2 Understanding Different Quantum Computers
Quantum computing is built on qubits as the fundamental units of quantum information. Broadly, qubit implementations can be categorized into two families: synthetic qubits, which are artificially engineered entities designed to effectively realize two-level quantum behavior, and natural qubits, which directly use the intrinsic quantum states of atoms, ions, or photons. Table 1 provides an overview of various key qubit technologies, representative commercial providers, and high-level performance characteristics.
Table 1: Quantum computing with different qubit technologies and their characteristics.
@@ 2 Understanding Different Quantum Computers
2.1 Quantum Computers with Synthetic Qubits
@@
2.2 Quantum Computers with Natural Qubits
@@
3 Review on Shor’s Implementations
3.1 Shor’s Algorithm
As discussed in [5], quantum computers excel at solving problems with modest classical input/output data but high computational complexity. Shor’s algorithm [1, 2] is a canonical example: it factors an $N$-bit integer in time polynomial in $N$, in stark contrast to the sub-exponential but super-polynomial complexity of the best-known classical factoring algorithms.
@@ 3.1 Shor’s Algorithm
3.2 Shor’s Implementation: State-of-the-Art
@@
In 2001, a group of IBM researchers presented the earliest experimental demonstration of Shor’s algorithm, successfully factoring $15=3\times 5$ using a seven-qubit liquid-state  [^1] Uses nuclear spins in molecules dissolved in a liquid as qubits. nuclear magnetic resonance (NMR) quantum computer [6]. While liquid-state NMR quantum computers are not considered a scalable route for solving larger problems, the experiment provided an important proof of principle, demonstrating the algorithm’s feasibility.
@@
Subsequent experiments implemented compiled versions of Shor’s algorithm using photonic qubits, again targeting the factorization of 15 [7, 8]. These demonstrations focused on generating and characterizing multi-qubit entanglement in circuits derived from Shor’s algorithm, but relied on heavy compilation that exploits prior knowledge of the factors.
@@
In 2012, factorization of 15 was successfully demonstrated on a superconducting quantum processor [9]. This milestone marked a significant advancement, showcasing the feasibility of implementing Shor’s algorithm on a quantum platform aligned with long-term scalability goals. In the same year, a photonic implmentation successfully factored 21 [10].
@@
In 2016, factorization of 15 was achieved using a trapped-ion implementation which used a qubit recycling technique [11]. Unlike previous compiled approaches, this demonstration did not rely on prior knowledge of the prime factors, and the design was argued to be scalable in principle, albeit still limited to a four-bit composite.
@@
More recent work was carried out on larger superconducting devices. In 2019, an attempt was made to factor 35 on the IBM Q System One [12]. However, the algorithm achieved only a 14% success rate due to accumulated noise, illustrating the fragility of deep circuits on current hardware. The latest IBM-based demonstation adopts a hybrid quantum-classical technique to variationally factor 253 (an 8 bit composite) [13]. While such methods are promising, they change the algorithmic structure and require careful interpretation when extrapolating to cryptographic scales.
@@
A number of works claim factorizations of larger integers using Shor-like circuits. As emphasized in the aptly-titled manuscript “Oversimplifying Quantum Factoring” [14], many of these experiments rely on heavily compiled circuits that incorporate knowledge of the answer, drastically lowering resource requirements, and sometimes reducing the problem to something closer to classical coin flipping than genuine quantum factoring. This underscores the importance of benchmarking implementations that preserve the true complexity of order-finding.
@@
In parallel, classical simulations of Shor’s algorithm on large-scale high-performance computing systems have advanced significantly. Recent work [15] reached 39 bits RSA key length on GPU clusters, validating the algorithmic structure and resource trends in a controlled environment. Nevertheless, the current (public) record is still held by classical algorithms such as the “Generalized Number Field-Sieve” (GNFS) which has reached 829 bits of key length [16].
@@
On the algorithmic side, the past two years have seen the first notable developments in the 30 years since Shor’s discovery in 1994. Recent work has lowered qubit counts [17, 18, 19] and improved asymptotic runtime [20], particularly by leveraging advances in quantum error correction and more efficient order-finding techniques. An important point to observe is the growing efficiency of incorporating progress in quantum error correction, which can be considered as the software layer between hardware (noisy physical qubits) and application (factoring). For example, updated resource estimates for factoring RSA-2048 have improved from around 20 million qubits and 8 hours of execution time [21] to fewer than one million qubits and roughly 5 days [22], under realistic assumptions about error rates and QEC overhead. Architectures that combine novel qubit designs with multimode memories could, at least in principle, reduce the required qubit count by another two orders of magnitude, down to approximately 13 436 qubits [23]. These developments emphasize that both hardware and algorithms are progressing and that resource estimates continue to evolve.
4 Shor’s Implementation on Existing Quantum Platforms
4.1 Setup and Notation
@@
However, with modern post-processing (Ekerå et al. [24]), the required phase precision can be cut to roughly $t\approx n+O(\log n)$ while maintaining high single-run success via limited classical searches, thereby reducing depth and repetitions [24, 25]. Thus, we allow ourselves to use a lower $t=10<12$ for the case $N=35$ while still considering the framework scalable  [^2] Further details on the experimental certification method and refinements regarding scalability in the context of Shor’s algorithm are provided in the Appendix A.1..
4.2 Public Quantum Platforms
@@
Table 2: Specifications of the ibm_torino Quantum Processor
@@
Table 2 summarizes the characteristics of ibm_torino, the backend used in our experiments. Two coherence-time metrics are listed. $T_{1}$ (energy relaxation time) measures how long a qubit prepared in the excited state ($|{1}\rangle$) remains there before decaying to the ground state ($|{0}\rangle$). Longer $T_{1}$ values indicate reduced energy relaxation and better stability. $T_{2}$ (dephasing time) quantifies how long a qubit maintains phase coherence in a superposition state; it captures the rate at which quantum information is lost due to environmental noise and control imperfections. While $T_{2}$ is upper-bounded by $2\,T_{1}$, it is typically significantly shorter in practice due to imperfect control and environmental noise.
@@
Our implementation is based on standard constructions for reversible arithmetic. We employ a Cuccaro-style ripple-carry adder [26] and a two’s-complement overflow/comparator method for modular addition [27], avoiding QFT-space adders [28]. Controlled powers of $a$ are implemented via precomputed exponents $a^{2^{k}}\bmod N$ within a textbook QPE order-finding scaffold [29, 30] and we target Qiskit Runtime Sampler V2 [31].
@@
We implement the parallel (non-iterative) phase-estimation variant of Shor’s order-finding rather than Kitaev’s iterative version [29]. Both are algorithmically equivalent: the dominant cost is the controlled modular exponentiation $U_{a}$, which scales as $\tilde{O}(n^{3})$ gates for $n=\lceil\log_{2}N\rceil$ with standard reversible adders and modular reduction, while the inverse QFT contributes $O(t^{2})$ gates for $t$ phase bits. The choice is therefore an engineering trade-off. Parallel QPE uses $t$ control qubits and a single inverse QFT, yielding a full $t$-bit outcome per shot. Iterative QPE reuses one control qubit over $t$ rounds with mid-circuit measurement and classical feed-forward, thereby reducing qubit count but increasing depth adding additional SPAM/latency errors. In our experiments, $t\approx 2\lceil\log_{2}N\rceil$ is modest, so the extra control qubits are acceptable and let us (i) avoid real-time feedback, (ii) exploit vectorized sampling, and (iii) apply a uniform multi-peak histogram analysis pipeline based on acceptance windows and binomial tests. In a fault-tolerant setting, either variant remains viable; our choice is driven by hardware constraints, not complexity.
@@ 4.3 Results (N=15, 21, 35)
We then executed the same circuits on ibm_torino, after transpilation to the device topology. For each modulus $N$, we focus on detecting a statistically significant excess of probability mass in the theoretically predicted windows around $sL/r$ relative to the uniform baseline. The formulas of the uniform baseline and the empirical probability used for the experiments, rigorously derived in the Appendix A.1, are the following:
@@
We consider three composite numbers: $N=15$, $N=21$, and $N=35$. In the Appendix subsection A.7, we further discuss the shape of the empirical histograms compared to the ideal noiseless expected distributions.
@@ 5 Limitations and Challenges
To achieve genericity at fixed width, the arithmetic would need to treat $N$ and related parameters as data rather than as compile-time structure. This would entail allocating an $n$-qubit modulus register $\mathrm{NREG}$ (and addend/multiplier registers as needed), replace constant-injection adders with register–register Cuccaro adders plus a reversible subtract-and-borrow comparator against $\mathrm{NREG}$, and implement modular multiplication as a variable shift-and-add with reductions modulo $\mathrm{NREG}$ (optionally via parameterized Fourier-space adders) [32]. Consequently, a comprehensive benchmark should employ $N$-agnostic arithmetic register–register add/compare/multiply mod $N$ or parameterized Fourier-space rotations so that a single width-fixed topology works for any $N<2^{n}$. Per-$N$ specialized circuits primarily demonstrate the behavior of bespoke instances and do not fully probe a device’s readiness for the generic operations required at cryptographically relevant key sizes.
@@
Quantum machine fidelity captures how accurately a device implements its nominal operations. Other factors such as qubit connectivity, coherence times, and readout error rates provide further metrics to estimate whether reliable results are produced when running Shor’s algorithm on real hardware. While we had access to calibration data for ibm_torino (Table 2), these values represent median performance and can vary substantially over time. In practice, we observed non-negligible day-to-day fluctuations that occasionally degraded circuit behavior to the point of rendering runs effectively unusable. As a result, some experiments had to be discarded as outliers corresponding to unfavorably calibrated device states. Below are presented the statistics of some error metrics over the months preceding and succeeding the experiments, for qubit q_0. A more systematic characterization of worst-case behavior-for instance, via repeated sampling over many calibration cycles-would be necessary to rigorously assess reliability guarantees. However, such an investigation would require significantly more QPU access time and lies beyond the scope of this study.
Table 3: Calibration statistics for ibm_torino machine over the months preceding and succeeding the experiments, for qubit $q_{0}$
@@
Our experiments focus primarily on detecting a statistically significant quantum signal in QPE histograms, rather than fully establishing polylogarithmic scaling of the expected runtime for a fixed acceptance parameter $k$ (see Appendix A.6). For each small modulus $N$, we tune window sizes and shot counts to test whether the mass in acceptance regions exceeds a uniform baseline. This is sufficient to detect coherence in the presence of noise but does not yet constitute a full scalability study.
@@
6 Trends in Quantum Computing Developments
Investment in quantum technologies has surged in recent years. Global public and private funding has been estimated to exceed \$55 billion [^7] https://www.qureca.com/quantum-initiatives-worldwide/, spanning quantum computing, communication, sensing, and related areas. Many governments have launched national or regional initiatives, such as the U.S. National Quantum Initiative [^8] https://www.quantum.gov/, the European Quantum Flagship  [^9] https://qt.eu/ and major programs in China and other countries, aimed at securing leadership in this emerging and potentially transformative field [33]. In parallel, private sector funding has grown rapidly, with large technology companies such as IBM, Google, Microsoft, and Amazon heavily investing in quantum computing, alongside startups such as IonQ, Quantinuum, Rigetti, PsiQuantum, Xanadu, and others, which are developing specialized quantum systems. Table 4 illustrates selected quantum computing roadmaps announced by industrial players, grouped by qubit technology.